If the perimeter of a triangle with a 90 degree angle is 1400, and the hypotenuse is 600, what are the other 2 sides?
Answer by stanbon(75887)

If the perimeter of a triangle with a 90 degree angle is 1400, and the hypotenuse is 600, what are the other 2 sides?
--------------------
Sum of the other 2 sides = 800
-----
You might try this.
x+y = 800
x^2 + y^2 = 1400^2
----
Solve for x and for y.
==========================
Cheers,
Stan H.